# UnMarker: A Universal Attack on Defensive Watermarking

### v1

@@ A.4. Baseline Results Against Naive Image Manipulations
Figure 4. Otputs of the three removal attacks. The VAEAttack over-smooths images, resulting in them losing crucial information and occasionally looking cartoonish. Compared to the DiffusionAttack, UnMarker’s outputs better resemble the watermarked images. The DiffusionAttack also introduces semantic incoherences such as additional teeth in the case of PTW and omits identifying details such as freckles on the forehead for StegaStamp. PTW’s watermarked image itself is of inferior quality with an unnatural patch. The DiffusionAttack retains this artifact while UnMarker eliminates it, enhancing the image.
A.5. Effects of Adversarial Filtering
@@
Table 6. UnMarker’s performance against semantic schemes with and without using optimizable filters.
@@
without filters     30.45 0.04 93.13% 100% 49.59 0.04 0.0147 86%
Despite our thorough analysis substantiating the choice to perform filtering-based optimizations to defeat low-frequency watermarks, this attack has two major changes compared to our high-frequency disruptive procedure. While one introduces the filtering-based approach, the other replaces the loss function that is obviously unsuitable for low-frequency distortions, as explained in §5.2.
The reader may wonder whether replacing the loss function alone can achieve the same outcome, eliminating the need for computationally expensive filters. Yet, based on our arguments explaining how direct input optimizations can only be of limited applicability against low-frequency watermarks due to their non-systemic nature, this is not the case. To validate this claim, we compare UnMarker against the two semantic (low-frequency) schemes, TRW and StegaStamp, with and without our filters. In both experiments, the loss is the same low-frequency function from §5.2.4.
The results in Table 6 unequivocally prove the filters necessary. While TRW, which the previous experiments prove more fragile, experiences a slight drop in performance without the filters, including them is far more effective, bringing its detection rate to $40\%$ (compared to $86\%$ without filters). StegaStamp remains completely unaffected when the filters are excluded, retaining a $100\%$ detection rate. With our filters, this drops to $43\%$. We also verified that the inability of the filter-free attack to break these schemes persists even if the perturbation budget slightly increases. Note that without filters, we obtain slightly better similarity scores since the filters encourage necessary structural changes. Yet, even with the filters, these scores still prove that UnMarker retains quality.
A.6. Outputs of UnMarker vs. The SOTA
Fig. 4 displays examples of watermarked images and the outputs obtained by subjecting them to advanced removal attacks.

### v2

@@ Appendix C Baseline Results Against Image Manipulations
Figure 4: Outputs of the three removal attacks. The VAEAttack over-smooths images, resulting in them losing crucial information and occasionally looking cartoonish. Compared to the DiffusionAttack, UnMarker’s outputs better resemble the watermarked images. The DiffusionAttack also introduces semantic incoherences such as additional teeth in the case of PTW and omits identifying details such as freckles on the forehead for StegaStamp. PTW’s watermarked image itself is of inferior quality with an unnatural patch. The DiffusionAttack retains this artifact while UnMarker eliminates it, enhancing the image.
The kernel size for Gaussian (Guided) Blurring is $5{\times}5$. We exploit Guided Blurring’s edge-preserving properties to increase the std. (to $5.0$). To control the modulation at edges, these accept an additional color kernel’s std. (see §5.2), for which we have three values: 0.1, 1.00.11.00.1,\;1.00.1 , 1.0, and $5.0$. Yet, $1.0$ and $5.0$ are overly destructive and do not preserve quality, chosen only to explore robustness. All parameters are in Table IV.
Results: Most systems are robust to all practical manipulations. Despite resisting all other modifications, Yu2 breaks completely under cropping. Hence, unlike other schemes, it highly depends on spatial attributes [^3] Yu et al. [22] only evaluate a single ProGAN [55] architecture under manipulations, whereas we chose their StyleGAN2 [63] model.. Yu1, HiDDeN and TRW exhibit a considerable drop in the bit accuracies (inverse distance) after cropping, although their overall detection is not affected. This is not surprising, as the watermarking spectral modifications must be spread throughout the image to minimize visual cues, implying that a portion may be allocated to the background. Yet, they retain robustness by restricting most of these changes to central regions.
Yu1, StegaStamp and TRW resist all manipulations, while StableSignature can withstand realistic, non-obtrusive modifications, making it also robust for practical purposes. PTW is slightly less powerful, losing some of its efficacy under even non-destructive operations, such as Quantization with a factor of $10$ and JPEG with quality $80$ that decrease its performance to $46\%$ and $76\%$, respectively. HiDDeN exhibits similar behavior, resisting all realistic manipulations aside from JPEG that brings its detection to $43\%$ even when used with a reasonably good quality of $110$ or Blurring with $std.=1.0$ that lowers its detection score to $5.0\%$.
It is expected that some non-semantic systems (e.g., PTW, HiDDeN and StableSignature) begin to break under spectral manipulations such as JPEG and Blurring. As explained in §4.1, constructing spectral non-semantic watermarks should be done in an input-specific manner that is also persistent, which is a challenging task that will be restricted to volatile watermarks unless the scheme is sufficiently advanced. This explains why the more sophisticated StableSignature and PTW outperform HiDDeN. Yet, allowing for severe manipulations via stronger (Guided) Blurring breaks StableSignature as well, although this is only included to demonstrate this phenomenon that attests to the spectral nature of these watermarks and is not a valid attack (see §6.1). Quantization can cause sharp pixel shifts (high-frequency noise), explaining PTW’s vulnerability.
With few exceptions, which we attribute to a slight mismatch in the experimental settings, our findings align with those attained by the authors of the schemes that evaluate the same attacks with similar parameters.
Appendix D Effects of Adversarial Filtering
@@
TABLE VI: UnMarker’s performance against semantic schemes with and without optimizable filters.
@@
w/o filters  30.45 0.04 93.13% 100% 49.59 0.04 0.0147 86%
The reader may wonder whether the low-frequency loss alone can defeat semantic schemes, eliminating the need for our filters. Based on our arguments explaining how direct optimizations are of limited applicability against low-frequency watermarks due to their non-systemic nature (see §5.2), this is not the case. For validation, we compare UnMarker against semantic schemes with and without our filters. In both experiments, the loss is the same from §5.2.4. The results in Table VI prove the filters necessary. While TRW experiences a slight degradation without filters, including them is far more effective, bringing its detection to $40\%$ (compared to $86\%$ otherwise). StegaStamp is unaffected when filters are excluded, retaining $100\%$ detection. With filters, this drops to $43\%$. We also verified the filter-free attack’s ineffectiveness persists even if the perturbation slightly increases. Although similarity is slightly worsened by our filters as they encourage structural changes, the scores show that, even with filtering, UnMarker retains quality.
Appendix E Outputs of UnMarker vs. The SOTA
Fig. 4 displays examples of watermarked images and the outputs obtained by subjecting them to advanced attacks.
Appendix F Meta-Review
The following meta-review was prepared by the program committee for the 2025 IEEE Symposium on Security and Privacy (S&P) as part of the review process as detailed in the call for papers.
F.1 Summary
This paper introduces UnMarker, a universal attack method designed to compromise defensive image watermarking mechanisms.
F.2 Scientific Contributions
  • Identifies an Impactful Vulnerability
  • Provides a Valuable Step Forward in an Established Field
  • Independent Confirmation of Important Results with Limited Prior Research
F.3 Reasons for Acceptance
  1. 1.
    UnMarker demonstrates effectiveness against multiple watermarking schemes.
  2. 2.
    The paper provides valuable insights into the vulnerabilities of current watermarking methods by leveraging spectral amplitude analysis.